Sally can paint a room in 4 hours while it takes Steve 6 hours to paint the same room. How long would it take them to paint the room if they worked together?
Answer by stanbon(75887)

Sally can paint a room in 4 hours while it takes Steve 6 hours to paint the same room. How long would it take them to paint the room if they worked together?
--------------------------------------
Sally rate = 1/4 job/hr
Steve rate = 1/6 job/hr
-----------------
Together rate = 1/x
------------------------------
Equation:
1/4 + 1/6 = 1/x
Multiply thru by 12x
3x + 2x = 12
5x = 12
x = 12/5 = 2 2/5 hrs = 2hr 24 minutes
=========================================
Cheers,
Stan H.